OpenStudy (anonymous):

Will fan and give medal!!!! What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y = x +

OpenStudy (anonymous):

@Secret-Ninja @sleepyjess @GingerSnap1299 @OrangeCaramell

OpenStudy (campbell_st):

the slope intercept form is \[y = mx + b\] m is the slope, which you have been given b is the y- intercept. to find b, use the information from the point x= 6 and y = 2 plug those ito your equation to dins b hope it helps

OpenStudy (anonymous):

y - y1 = m (x -x1)

OpenStudy (anonymous):

is it fine if you answer some more questions? @campbell_st

OpenStudy (anonymous):

@campbell_st